Fama’s Llamas has a weighted average cost of capital of 8.9 percent. The company’s cost of equity is 13 percent, and its pretax cost of debt is 6.2 percent. The tax rate is 23 percent. What is the company’s target debt-equity ratio?

Please explain each step as I need to understand this conceptually.
Answer #1

After-tax cost of debt=6.2*(1-tax rate)

=6.2*(1-0.23)=4.774%

Let debt be $x

Equity be $y

Total=$(x+y)

WACC=Respective cost*Respective weight

8.9=(x/(x+y)*4.774)+(y/(x+y)*13)

8.9*(x+y)=4.774x+13y

8.9x+8.9y=4.774x+13y

x=(13-8.9)y/(8.9-4.774)

=0.9937y

Hence debt-equity ratio=debt/equity

=0.9937(Approx).